OpenStudy (anonymous):

Prove the following inequality when x is positive: 4x^10+x^8+4x^2+1≥2x^9+4x^5+2x

OpenStudy (anonymous):

Take the log of both sides

OpenStudy (anonymous):

log base x Log(4)+10+8+Log(4)+2+1 >= Log(2) + 9 + Log(4) + 5 + Log(2) + 1 Use properties of logs to combine and subtract from both sides Log(16) + 21 >= Log(16) + 15 21 >= 15

OpenStudy (anonymous):

I am not sure of you got to the first line. Won't it remain as log(base x)(4x^10+x^8+4x^2+1)≥log(base x)(2x^9+4x^5+2x) ? How could you break it up like that? Please explain. Thanks

OpenStudy (anonymous):

You're right I did it wrong, let me look for another way

OpenStudy (anonymous):

I couldn't find any better proof but, divide both sides by the right hand side to get (4x^10+x^8+4x^2+1)/(2x^9+4x^5+2x) > 1. Find the minimum of the left hand side by taking the derivative and solving for 0. (Ugly). You get x=.930187 for a minimum. (only positive minimum) Plugging that into the left hand side you get 1.224, which is always > 1.